Overweight baggage. Suppose weights of the checked baggage of airline passengers follow a nearly normal distribution with mean 40 pounds and standard deviation 2.5 pounds. Most airlines charge a fee for baggage that weigh in excess of 50 pounds. Determine what percent of airline passengers incur this fee. Round your answer to four decimal places. %
Answer #1

Solution :

Given that ,

mean = = 40

standard deviation = = 2.5

P(x > 50 ) = 1 - P(x < 50)

= 1 - P[(x - ) / < (50-40) /2.5 ]

= 1 - P(z <4 )

= 1- 1 = 0.0000

ans = 0.00%